Find the number c guaranteed by the mean value theorem for derivatives for f(x)=(x+1)^3, [-1,1]
1
Expert's answer
2019-05-15T14:17:26-0400

Solution. Using mean value theorem: Suppose f(x) is a function that satisfies all of the following.

1) f(x) is continuous on the closed interval [a;b].

2) f(x) is differentiable on the open interval (a;b).

Then there is a number c such that a<c<b and


"f'(c)=\frac {f(b)-f(a)} {b-a}"


Function  f(x)=(x+1)^3 is continuous on the closed interval [-1,1] and f(x) is differentiable on the open interval (-1,1).


"f(-1)=(-1+1)^3=0"

"f(1)=(1+1)^3=8"

Therefore

"f'(c)=\frac {8-0} {1-(-1)}=4"


Find the first derivative of the function f(x).


"f'(x)=3(x+1)^2"

As result get equation


"3(x+1)^2=4"

Find the roots of the equation.


"x+1=\frac {2} {\sqrt {3}} \Longrightarrow x_1=-1+\frac {2} {\sqrt {3}}"

"x_1 \isin (-1,1)"

"x+1=-\frac {2} {\sqrt {3}} \Longrightarrow x_2=-1-\frac {2} {\sqrt {3}}"

"x_2 \notin (-1,1)"


Answer.


"-1+\frac {2} {\sqrt {3}}"